# A Volumetric Approach to Privacy of Dynamical Systems

### v3

@@ Abstract
Information-theoretic metrics, such as mutual information, have been widely used to evaluate privacy leakage in dynamic systems. However, these approaches are typically limited to stochastic systems and face computational challenges. In this paper, we introduce a novel volumetric framework for analyzing privacy in systems affected by unknown but bounded noise. Our model considers a dynamic system comprising public and private states, where an observation set of the public state is released. An adversary utilizes the observed public state to infer an uncertainty set of the private state, referred to as the inference attack. We define the evolution dynamics of these inference attacks and quantify the privacy level of the private state using the volume of its uncertainty sets. We then develop an approximate computation method leveraging interval analysis to compute the private state set. We investigate the properties of the proposed volumetric privacy measure and demonstrate that it is bounded by the information gain derived from the observation set. Furthermore, we propose an optimization approach to designing privacy filter using randomization and linear programming based on the proposed privacy measure. The effectiveness of the optimal privacy filter design is evaluated through a production-inventory case study, illustrating its robustness against inference attacks and its superiority compared to a truncated Gaussian mechanism.
@@
Volumetric privacy measure; privacy protection; interval analysis; truncated Gaussian mechanism.
@@ 1.1 Motivation
Data sharing plays a pivotal role in enabling cooperative decision-making and optimization in dynamic processes. However, the exposure of such data may inadvertently reveal sensitive information. Specifically, correlations between shared metrics and underlying operational parameters can be exploited by adversaries to develop competitive and malicious strageties. This challenge highlights the critical need for methodologies that preserve data utility while ensuring rigorous privacy protection for dynamic systems.
Information-theoretic metrics, including mutual information, directed information, and conditional entropy, have been extended to quantify privacy leakage in dynamic systems. Despite their utility, these metrics have notable limitations. For instance, as highlighted in [8, 30], mutual information evaluates privacy leakage in an average sense, which can overlook privacy breaches for infrequent realizations of private variables. The computational complexity of these measures is another challenge, as they require evaluating leakage for all possible realizations of random variables [25, 32, 11]. Most existing information-theoretic approaches are focused on stochastic systems and typically assume complete knowledge of the probability distributions of system states. However, this assumption does not hold in systems that are influenced by unknown but bounded (UBB) noise.
To address these challenges, this paper introduces a volumetric privacy measure based on the uncertainty set of the private variable. The proposed approach is applicable to both deterministic and stochastic systems and eliminates the need for prior knowledge of probability distributions.
@@ 1.2 Contributions
This paper addresses the privacy protection problem for dynamic systems with UBB noise, as illustrated in Fig. 1. In this framework, the system state is partitioned into two categories: the public state $X_{k}$ and the private state $Y_{k}$, both of which belong to a known bounded set. The observation of the system states is represented by the set $\mathcal{M}_{k|k}^{x}$ which contains the actual public state $X_{k}=x_{k}$. An untrusted third party, i.e., the adversary, uses $\mathcal{M}_{k|k}^{x}$ to infer the private state by constructing an uncertainty set $\mathcal{Y}_{k|k}$, which is referred to as the inference attack.
Figure 1: The inference attack.
This system setup is inspired by the production-inventory problem discussed in Section 2. In such scenarios, inventory information can aid distributors in optimizing sales. However, it also enables adversaries to infer sensitive information, such as the production rate, which reveals a company’s private data, including production efficiency, strategies, and supply chain dynamics.
The contributions of this paper are summarized as follows:
  • Definition of Inference Attack: We define the adversary’s inference attack using set operations.
  • Volumetric Privacy Measure: A novel privacy measure is proposed based on the volume of the uncertainty set for the private state.
  • Privacy Level Computation: We develop a computational method utilizing interval analysis to evaluate the privacy level.
  • Optimal Privacy Filter Design: An optimization-based approach is introduced to design privacy filters, based on randomization and linear programming.
Additionally, we validate our approach using the production-inventory example. The results demonstrate that the adversary can infer sensitive production rates, through inventory data. However, this leakage can be significantly mitigated under the proposed optimal privacy filter design. Also, compared with the truncated Gaussian mechanism, our privacy filter design achieves lower data distortion while maintaining the same privacy level.
1.3 Related work
Differential privacy was initially defined to hide the presence of an individual or a specific record in the responses of queries for static datasets. The authors in [15] extended the differential privacy to dynamic systems, and proposed the differentially private Kalman filter for linear systems. In [5], the authors proposed a general architecture for private filtering by approximating the desire filter in a differential private way, and studied the application of differential privacy in distributed control and optimization. In [21], the authors studied the differential privacy preserving average consensus problem to protect the initial state in distributed control by adding and abstracting random noises. The authors in [12] connected the input observability with differential privacy and proposed the privacy preserving controller design method. The authors in [33] studied the minimal amount noises added to multi-agent systems for differential privacy based on the minimal observability subspace. They proposed node-based and edge-based privacy-preserving mechanisms based on the optimization of the minimum added noise, and discussed the trade-offs between utility and privacy. In [34], the authors proposed several cost-friendly differential privacy-preserving schemes by randomly adjusting the charge-discharge rate of a batter based on a modified Laplace distribution.
Information-theoretic protection methods use conditional entropy, mutual information, etc, from information theory, to measure privacy leakage. In [22], the authors studied the privacy filter design for a Markov chain to hide the private state when transmitting the monitored state to a receiver, which was approximately solved with a greedy algorithm based on the convex optimization. The authors in [3] designed a private filter to protect the state of a hidden Markov chain via minimizing the mutual information with the constraint of the observation utility. In [31, 24], the authors studied the privacy-aware filter design for linear systems, using the directed information between the system state and the output of the filter as the privacy measure. It was proved that the optimal filter can be realized with a Kalman filter and an additive Gaussian mechanism.
The authors in [32] studied the structural properties of the privacy-aware state estimation problem based on the dynamic programming decomposition. They approximated the privacy leakage with variational techniques and solved the estimation problem via policy gradient approaches. In [23], the conditional entropy of the private state given measurements and controls was used as a penalty term in the partially observable Markov decision process (POMDP) optimization, to avoid the adversary accurately estimate the system state. They showed that the state obfuscation problem is a standard POMDP optimization problem in which the cost-to-go function is concave in the belief state. In [16], the authors considered to protect the privacy of the electricity demand of a household to avoid the adversary infers the indoor activity. They utilized the rechargeable battery to hide the actual electricity demand and obtained the optimal charging policy via minimizing the privacy leakage measured by mutual information. However, the classical information-theoretic metrics for privacy might have some limitations when facing one-try attack as discussed in [30], which motivates the development of $R\acute{e}nyi$ min-entropy [30], maximal leakage [9] and pointwise maximal leakage [28], etc.
In most existing privacy protection literature for dynamic systems, it is assumed that the system are disturbed with unbounded noise that has probabilistic properties, e.g., Gaussian distribution. However, there are also dynamic systems are driven with UBB noise without probability assumptions. The privacy concept for these systems are not well developed. The authors in [6, 7] designed differentially private set-based estimator to protect the privacy of measurements based on a truncated noise distribution. In [13], the authors proposed the guaranteed privacy extent from differential privacy to hide the measurement sensor identities, and provided the optimization method for the $\mathcal{H}_{\infty}$ optimal privacy preserving interval observer design. Furthermore, in [26, 27], authors studied the state opacity problems where the discrete state space is divided into disjoint sets of secrete and non-secrete states, and the system outputs from secrete and non-secrete sets are indistinguishable. The authors in [19] proposed $\delta-$approximate initial-state opacity as a secrete notion for discrete-time systems with continuous states, and developed verification approaches based on Barrier certificates.
The proposed volumetric privacy measure in this work is inspired by set-membership estimation methods for systems with UBB noise [20]. Most set-membership methods represent states as bounded geometric sets, such as intervals [10], zonotopes [14], or ellipses [4], where the volumes of these sets correspond to estimation uncertainty. Building on these methods, we consider the adversary’s ability to infer the private state set based on observations of the public state. Consequently, a larger volume of the uncertainty set implies a broader range of potential private states that could be consistent with the observed public state set. This motivates our use of the volume of the private state set as a privacy measure.
This work differs from prior studies, such as [6], [7], and [13], which focus on ensuring the differential privacy of system states. In contrast, we address the problem of protecting private states while publicly releasing non-sensitive states, as illustrated in Fig. 1. Our approach uses a volumetric privacy measure, which also distinguishes it from state opacity problems [26], [27], and [19], where the state space is partitioned into secret and non-secret sets.
Furthermore, state opacity problems focus on guaranteeing privacy by ensuring that at least one non-secret state produces an observation indistinguishable from the secret state. In these frameworks, enlarging the volume of the non-secret state set does not enhance privacy, as privacy is determined solely by the existence of indistinguishable outputs, not by the size of the state set. Differently, within our framework, a larger private state set is preferred, as it increases the adversary’s uncertainty in inferring the private state.
@@ 1.4 Outline
The rest of the paper is organized as follows. Section 2 introduces the system model and the inference attack. Section 3 defines the volumetric privacy measure, provides computational approaches for privacy level evaluation, and discusses the properties of the proposed measure. Section 4 presents an optimal privacy filter design to mitigate privacy leakage while maintaining a certain utility level. Section 5 presents numerical results, followed by the conclusions in Section 6.
@@ 1.5 Notation
We use italic letters to denote the set of unknown variables, e.g., $\mathcal{X}$ and $\mathcal{Y}$ for $X$ and $Y$. We use the vector $\left[\begin{array}[]{c}\underline{X}\\ \overline{X}\\ \end{array}\right]$ to describe the interval $\mathcal{X}$, i.e., $\mathcal{X}=\left\{X|\underline{X}\leqslant X\leqslant\overline{X}\right\}$, where $\underline{X}$ and $\overline{X}$ are the lower and upper endpoints, respectively. The interval $\mathcal{X}$ can also be represented with
$\mathcal{X}=\left\{c^{x}+\mathrm{diag}\left(p^{x}\right)\alpha:\alpha\in\mathcal{R}^{n_{x}},\left|\alpha\right|_{\infty}\leqslant 1\right\},$
where $c^{x}=\frac{\overline{X}+\underline{X}}{2}$ and $p^{x}=\frac{\overline{X}-\underline{X}}{2}$ are the center point and radius of the interval. Besides, given the block matrix $A=\left[A_{1},A_{2}\right]$, the multiplication between the matrix $A$ and the interval $\mathcal{X}$ is defined as $A\mathcal{X}=A_{1}\underline{X}+A_{2}\overline{X}$. Also, if $\mathcal{X}$ and $\mathcal{Y}$ are intervals, then we use $\mathcal{X}+\mathcal{Y}$ to denote the interval $\left[\begin{array}[]{c}\underline{X}+\underline{Y}\\ \overline{X}+\overline{Y}\\ \end{array}\right]$, and use $\mathcal{X}-\mathcal{Y}$ to represent the interval $\left[\begin{array}[]{c}\underline{X}-\overline{Y}\\ \overline{X}-\underline{Y}\\ \end{array}\right]$. As for the non-interval set $\mathcal{Z}$, we use $A\mathcal{Z}$ to denote the new set $\left\{AZ|Z\in\mathcal{Z}\right\}$, and use $\mathcal{Z}\oplus\mathcal{R}$ to represent $\left\{Z+R|X\in\mathcal{Z},R\in\mathcal{R}\right\}$. Furthermore, the 1-norm of the column vector $b$ with n dimensions is defined as $\left\|b\right\|_{1}=\sum_{i=1}^{n}\left|b\left(i\right)\right|$, and $b^{\top}$ is the transpose of $b$. The vector $\mathbf{1}_{n_{x}}$ denotes a column vector of ones with $n_{x}$ dimensions, while $I_{n_{x}\times n_{x}}$ represents an identity matrix of size $n_{x}\times n_{x}$.
@@ 2.1 System Model
where $A_{1}$ and $A_{2}$ are invertible, $Y_{k}\in\mathcal{R}^{n}$ is the private state, $X_{k}\in\mathcal{R}^{n}$ is the public state to be released, $W_{k}^{x}\in\mathcal{W}_{k}^{x}\subseteq\mathcal{R}^{m}$ and $W_{k}^{y}\in\mathcal{W}_{k}^{y}\subseteq\mathcal{R}^{m}$ are the UBB noises. Also, the initial public and private states belong to $\mathcal{X}_{0|-1}$ and $\mathcal{Y}_{0|-1}$, respectively. When $A_{3}$ is zero, the model $\mathbf{G_{1}}$ can be regarded as a dynamic system with the public state $X_{k}$ and the private input $Y_{k}$.
For analysis convenience, we assume that $\mathcal{X}_{0|-1}$, $\mathcal{Y}_{0|-1}$, $\mathcal{W}_{k}^{x}$ and $\mathcal{W}_{k}^{y}$ are intervals. For other types of bounded sets, such as zonotopes [14] and ellipses [4], the analysis is left for future investigation. Note that we do not assume the prior knowledge of probability distributions of the process noises, e.g., Gaussian distribution, which is different from most existing literature in privacy protection for dynamic systems. Furthermore, we assume that the adversary has full knowledge of system model $\mathbf{G_{1}}$ and will collect information of the public state to infer the private state.
2.2 Motivating Example
Consider the problem of privacy leakage related to the production rate in inventory control systems. To enhance sales, companies often need to share inventory levels with distributors. However, inventory levels are closely correlated with the production rate, which serves as a key indicator of the factory’s production strategy, operational efficiency, and supply chain dynamics. Therefore, the production rate constitutes highly sensitive information. If accessed by an adversary, this data could be exploited to develop competitive strategies, potentially compromising the company’s market position and operational integrity.
The inventory of production can be modeled as follows [17, 29],
$X_{k}=X_{k-1}+Y_{k-1}-W_{k}^{x},$ (4)
where $X_{k}$ is the inventory, $Y_{k}$ is the production rate, and $W_{k}^{x}$ is the uncertain demand. The production rate is usually chosen to maintain the inventory around a desired level, which can be modeled as follows,
$Y_{k}=A_{3}X_{k-1}+A_{4}Y_{k-1}+W_{k}^{y},$ (5)
where the uncertain process noise $W_{k}^{y}$ is due to the adversary’s limited knowledge about the company’s production capability.
When the uncertainty of demand and production rate is small, e.g., $W_{k}^{x}=D^{x}$ and $W_{k}^{y}=D^{y}$ are constant, then the adversary can approximately reconstruct the production rate with the inventory level via
$Y_{k-1}\approx X_{k}-X_{k-1}+D^{x},$ (6)
and
$Y_{k}\approx A_{3}X_{k-1}+A_{4}Y_{k-1}+D^{y},$ (7)
which causes privacy leakage and increases potential risks of business competition.
However, in practical scenarios, the uncertainty associated with demand and production capability is bounded but cannot be ignored. Inferring the private production rate, therefore, becomes a complex but still feasible task. Due to the presence of unknown noise terms that belong to bounded sets without probabilistic assumptions, multiple private states may correspond to the same public state with equal probability, leading to the formation of an uncertainty set. To address this challenge, we define the adversary’s inference attack on the private state using set operations in the following sections.

### v4

# A Volumetric Privacy Measure for Dynamical Systems With Bounded Disturbance

@@ Abstract
This paper presents a volumetric privacy framework for dynamical systems subject to bounded disturbances, developed without requiring prior knowledge of their probability distributions. We consider systems with both public and private states, where a set containing the public state is shared as the observation. An adversary is assumed to execute an inference attack by exploiting the observed public state set to estimate an uncertainty set for the private state. The volume of this inferred set quantifies the adversary’s estimation uncertainty and serves as the proposed volumetric privacy metric. Approximate set-membership estimation techniques are developed to compute the private-state uncertainty set, and the properties of the privacy measure are analyzed, demonstrating that it is bounded by the information gain from the observation set. Furthermore, an optimization-based privacy filter design problem is formulated, employing randomization and linear programming to enhance the volumetric privacy level. The effectiveness of the proposed approach is validated through a production–inventory case study. Results show that the optimal privacy filter significantly improves robustness against inference attacks and outperforms two baseline mechanisms based on additive noise and quantization.
@@
Volumetric privacy measure; privacy protection; interval analysis; bounded disturbance.
@@ 1.1 Motivation
Data sharing plays a pivotal role in enabling cooperative decision-making and optimization in dynamic processes. However, the exposure of such data may inadvertently reveal sensitive information. Specifically, correlations between shared metrics and underlying operational information can be exploited by adversaries to develop competitive and malicious strageties. This challenge highlights the critical need for methodologies that preserve data utility while ensuring privacy protection for dynamic systems.
Dynamical systems subject to bounded disturbances without knowledge of their underlying distributions provide a natural framework for modeling numerous practical applications involving sensitive information, e.g., inventory–production systems and connected vehicles. However, the notion of privacy in such systems remains insufficiently explored. Within the context of set-membership estimation, the states of these systems can be represented by geometric sets, such as ellipsoids or zonotopes, whose volumes quantify the degree of inference uncertainty.
Motivated by these considerations, this paper investigates the notion of volumetric privacy for systems affected by bounded disturbance. We develop privacy-preserving strategies aimed at maximizing an adversary’s uncertainty, i.e., the volume of uncertainty set, in inferring private states. The proposed approaches are applicable to both deterministic and stochastic systems, without requiring prior knowledge of the underlying probability distributions.
1.2 Related Work
Stochastic approaches to privacy primarily include differential privacy and information-theoretic methods. Differential privacy (DP) [1] has been incorporated into dynamic settings through differentially private Kalman filtering [2], DP-preserving average consensus via noise injection [3], and minimal-noise mechanisms for multi-agent systems based on observability properties [4]. Recent work [5] introduced a trace-based variance–expectation ratio to quantify topology preservation and derived optimal noise designs, while [6] provided a comprehensive survey of DP in dynamical systems. In parallel, information-theoretic approaches quantify privacy leakage using mutual information or conditional entropy. Recent studies include mutual-information-based private filtering for hidden Markov models [7], directed-information-based privacy filters for linear systems [8, 9], and recent extensions to partially observable Markov decision processes addressing privacy-aware estimation and control [10, 11].
Most existing studies focus on deterministic or stochastic systems with unbounded noise and known distributions, leaving privacy protection for systems subject to unknown-but-bounded disturbances relatively unexplored. Recent works [12, 13] developed differentially private set-based estimators using truncated noise, while [14] introduced guaranteed privacy concepts and optimization methods for $\mathcal{H}_{\infty}$-based privacy-preserving interval observers. State-opacity-based methods [15, 16] ensured indistinguishable outputs between secret and non-secret states but did not quantify the associated estimation uncertainty. Note that set-membership estimators typically characterize uncertainty through bounded geometric sets such as intervals [17], zonotopes [18], or ellipsoids [19], where the corresponding set volume naturally describes the amount of estimation uncertainty. Motivated by this, we analyze privacy leakage in systems with private and public states and propose a volumetric approach that maximizes the private-state set volume, thereby enhancing privacy while explicitly accounting for geometric effects under inference attacks.
There are some related deterministic approaches to privacy without adding noise, e.g., plausible deniability [20] and noiseless privacy [21]. In [20], privacy leakage in deterministic systems was measured by the volume of reachable state sets, and was determined by the observability. However, this framework does not apply to systems with bounded disturbance, where inference uncertainty depends on both observability and disturbance. Moreover, the problem of privacy filter design was not addressed in [20], whereas we propose a concrete design using randomization and optimization. The work in [22] addressed parameter privacy in deterministic systems via constrained convex generators (CCGs), which differs from our private state protection setting. Also, defense strategies in [22] involve ceasing information sharing or altering parameters, which might be unsuitable for fixed-parameter systems with continuous communication. As discussed in Sec. 3.1, the complexity of CCG-based inference grows exponentially with time, motivating our use of interval analysis for computational efficiency.
Noiseless privacy [21] and non-stochastic privacy [23] employed non-stochastic information-theoretic approaches to limit information leakage, assuming static private-variable domains and without accounting for temporal dependencies in sequential data. While noiseless privacy, non-stochastic privacy, and our volumetric privacy all achieve privacy through the release of bounded outputs. However, in our setup, dynamical systems subject to bounded disturbance have time-variant private-state reachable sets that can be recursively estimated, enabling dynamic leakage evaluation. Building on this insight, the proposed volumetric privacy filter dynamically evaluates the private state set and adapts the observation accordingly, thereby achieving higher privacy levels with lower data distortion, as shown in Sec. 5.
Finally, other deterministic privacy-preserving approaches often exploit observability reduction or state decomposition to protect private information in multi-agent systems. For example, the authors in [24] established a connection between network privacy and its observability space, proposing a privacy-aware communication protocol that achieves average consensus while protecting initial states. In [3], the trace of the observability Gramian was employed to quantify information leakage through an intruder node, and an online optimization strategy was proposed to adapt communications in order to minimize such leakage. Privacy-preserving consensus designs were also proposed in [25, 26], including augmented states and novel consensus algorithms with simultaneous accuracy, resilience, and privacy guarantees. State decomposition methods, as in [27], split each node’s state into randomized components to prevent disclosure of individual states during consensus.
@@ 1.3 Contributions
This paper studies privacy for dynamic systems with bounded disturbance as shown in Fig. 1, where the system state is split into public $X_{k}$ and private $Y_{k}$, and an adversary uses the public observation set $\mathcal{M}_{k|k}^{x}$ to infer $Y_{k}$ via an uncertainty set $\mathcal{Y}_{k|k}$.
Figure 1: The system setup.
The primary contribution of this work is the development of an extensible framework for privacy analysis and mitigation in dynamic systems subject to bounded disturbance. This contribution can be summarized in three principal aspects. (1) Volumetric Privacy Measure: We introduce a privacy metric based on the volume of the estimated private-state set obtained via set-membership estimation given the available observations. (2) Privacy Level Computation: We develop computational methods to quantify privacy level and prove that the relevant privacy leakage is bounded by the information gain from the observations. (3) Optimal Privacy Filter: Since the inappropriate choice of the observation set would lead to privacy leakage of the private state, we design a randomized, optimization‐based filter that perturbs and then refines observations to maximize inference uncertainty of attackers. Finally, the proposed framework is demonstrated on a production–inventory case study, showing that our privacy filter significantly reduces the adversary’s capability to estimate the private production rate.
@@ 1.4 Outline
The rest of the paper is organized as follows. Section 2 introduces the system model, the inference attack and defines the volumetric privacy and utility measure. Section 3 provides computational approaches for privacy level evaluation, and discusses the properties of the proposed measure. Section 4 presents an optimal privacy filter design to mitigate privacy leakage while maintaining a certain utility level. Section 5 presents numerical results, followed by the conclusions in Section 6.
@@ 1.5 Notation
We use italic letters to denote the set of unknown variables, e.g., $\mathcal{X}$ and $\mathcal{Y}$ for $X$ and $Y$. For the non-interval set $\mathcal{Z}$, we use $A\mathcal{Z}$ to denote the set $\left\{AZ|Z\in\mathcal{Z}\right\}$, and use $\mathcal{Z}\oplus\mathcal{R}$ to represent $\left\{Z\!+\!R|Z\in\mathcal{Z},R\in\mathcal{R}\right\}$. Furthermore, the 1-norm of the column vector $b$ with n dimensions is defined as $\left\|b\right\|_{1}=\sum_{i=1}^{n}\left|b\left(i\right)\right|$ with the absolute value $\left|b\left(i\right)\right|$, and $b^{\top}$ is the transpose of $b$. The 1-norm of the matrix $A$ is defined as $\left\|A\right\|_{1}\!\frac{\Delta}{=}\!\sum_{i,j}^{n,m}{\left|a_{i,j}\right|}$. The vector $\mathbf{1}_{n_{x}}$ denotes a column vector of ones with $n_{x}$ dimensions, while $I_{n_{x}\!\times\!n_{x}}$ represents an identity matrix of size $n_{x}\times n_{x}$. The operator $diag(v)$ denotes a diagonal matrix constructed from the vector $v$.
@@ 2.1 System Model
where $A_{1}$ and $A_{2}$ are invertible, $Y_{k}\in\mathcal{R}^{n}$ is the private state, $X_{k}\in\mathcal{R}^{n}$ is the public state to be released, $W_{k}^{x}\in\mathcal{W}_{k}^{x}\subseteq\mathcal{R}^{m}$ and $W_{k}^{y}\in\mathcal{W}_{k}^{y}\subseteq\mathcal{R}^{m}$ are the unknown disturbance with bounded sets $\mathcal{W}_{k}^{x}$ and $\mathcal{W}_{k}^{y}$. Note that the underlying probability distributions of $W_{k}^{x}$ and $W_{k}^{y}$ are unknown. Besides, the initial public and private states belong to $\mathcal{X}_{0|-1}$ and $\mathcal{Y}_{0|-1}$, respectively. We assume that the adversary has full knowledge of system model $\mathbf{G_{1}}$ and will collect information of the public state to infer the private state.
2.2 Motivating Examples
In this subsection, we consider two motivating examples to illustrate the necessity of protecting privacy of systems $\mathbf{G_{1}}$.
Production-inventory system: In supply chain management [28, 29], the inventory level $X_{k}$ and the production rate $Y_{k}$ evolves according to $\mathbf{G}_{1}$. While firms may disclose inventory information $X_{k}$ to distributors to boost sales, the production rate $Y_{k}$ contains sensitive strategic information such as production efficiency and supply chain operations. Since $X_{k}$ and $Y_{k}$ are correlated, releasing $X_{k}$ directly risks revealing private production details. Therefore, it is necessary to transform or mask observations to preserve the privacy of $Y_{k}$ while maintaining the utility of public inventory data $X_{k}$.
Traffic management system: In intelligent transportation, vehicles may report their velocities to a central controller to optimize traffic flow, e.g., by adjusting the speed limit on highways. Given bounded disturbances from environmental factors like uneven ground, the vehicle dynamics fit the model $\mathbf{G}_{1}$ with unknown-but-bounded disturbance. Here, velocity $X_{k}$ can be considered public data used for traffic management, while position $Y_{k}$ is private, as it can be used to identify individual vehicles. To protect location privacy, vehicles may intentionally report blurred or randomized velocity observations that preserve system utility but reduce the risk of precise location inference.